OpenStudy (anonymous):

Having trouble finding the correct substitutions & tricks for this one: \[\int\limits_{-1}^{1} \frac{7xe^{2x} }{(1 + 2x)^2} dx\] (don't need to evaluate, will give a medal)

OpenStudy (anonymous):

u= (exp(2x))/(1+2x)

OpenStudy (anonymous):

amazing, right? you get just what you want in one step. problem must have been cooked

OpenStudy (anonymous):

but be very careful with the integral, because it is improper as the integrand is undefined at \(x=-\frac{1}{2}\)

OpenStudy (anonymous):

very very careful!

OpenStudy (anonymous):

Right so it's just a simple u substitution and then splitting the integral up into a sum of two smaller intervals avoiding the -0.5, correct?

OpenStudy (anonymous):

u can use this \[\lim_{t \rightarrow -1/2}(\int\limits\limits_{-1}^{t}f(x)dx+\int\limits\limits_{t}^{1}f(x)dx)\]